III – Find the approximate p-value given the following information:



1. A two-tailed test; computed t-value = 1.845; df = 10




2. A left-tailed test; computed t-value = 2.295; df = 21




3. A right-tailed test; computed t-value = 1.565; df = 27




4. A two-tailed test; computed t-value = 2.452; df = 12




1
Expert's answer
2022-05-30T16:39:37-0400

1.

Two-tailed test: "p-value = 2 \cdot min\{cdf(x) , 1 - cdf(x)\}"


"p=0.094819"

2.

Left-tailed: "p-value = cdf(x)"


"p=0.983932"

3.

Right-tailed: "p-value = 1-cdf(x)"


"p=0.983932""p=0.064615"

4.

Two-tailed test: "p-value = 2 \cdot min\{cdf(x) , 1 - cdf(x)\}"


"p=0.030481"